### v1

@@ 1 Introduction
In this study, we propose an enhanced Trust Region Sequential Convex Optimization algorithm tailored specifically for trajectory planning of multi-drone thermal screening systems in complex urban environments. Our contributions are threefold: First, we introduce an improved convexification method that effectively addresses the nonlinear and non-convex constraints characteristic of urban drone operations, resulting in enhanced convergence stability and computational efficiency. Second, the integration of a dynamically adaptive trust-region strategy ensures robust trajectory refinement, significantly reducing computation time while maintaining path optimality and safety constraints. Finally, we validate the effectiveness and superiority of our algorithm through comprehensive simulations in realistic urban scenarios, demonstrating marked improvements in thermal screening coverage, trajectory smoothness, and overall system performance compared to conventional convex optimization methods. This research thus represents a significant advancement toward practical and scalable multi-drone thermal screening solutions in urban settings.
@@ 3.1 Introduction to the Trust Region Sequential Convex Optimization Algorithm
3.2 Enhancements for Thermal Screening Trajectory Planning
@@
A dynamically adaptive trust region strategy is integrated to balance local exploration and global convergence. The trust region constraint limits the state deviation from the reference trajectory $(x_{\text{ref}},y_{\text{ref}},z_{\text{ref}})$ during each iteration:
$\|x(t)-x_{\text{ref}}(t)\|_{2}\leq\delta_{x},$ (20)
$\|y(t)-y_{\text{ref}}(t)\|_{2}\leq\delta_{y},$
$\|z(t)-z_{\text{ref}}(t)\|_{2}\leq\delta_{z},$
where $\delta_{x},\delta_{y},\delta_{z}$ are adaptive radii adjusted based on the consistency between predicted and actual objective function improvements. The radii are updated based on the consistency ratio $\rho$ between the actual reduction in the objective function $J$ and the reduction predicted by the linearized model $J_{\text{pred}}$:
$\rho=\frac{J(\mathbf{x}^{\text{ref}})-J(\mathbf{x})}{J_{\text{pred}}(\mathbf{x}_{\text{ref}})-J_{\text{pred}}(\mathbf{x})},$ (21)
where $\mathbf{x}_{\text{ref}}$ is the reference trajectory and $\mathbf{x}$ is the solution of the current convex subproblem. Based on $\rho$, the trust region radii are updated as follows:
$\delta_{i}^{k+1}=\begin{cases}\min(\delta_{i}^{k},\delta_{\text{max}})&\text{if }\rho>\rho_{0},\\ \delta_{i}^{k}&\text{if }\rho\leq\rho_{0}\par\end{cases}$ (22)
where $\rho_{0}$ is a threshold parameter in the trust region radius adjustment strategy and $\delta_{\text{max}}$ is the maximum allowable radius. This prevents excessive linearization errors, ensuring stable convergence even in cluttered environments.
@@ 3.3 Mathematical Formulation and Optimization Approach
$\min$ $J=J_{\text{smooth}}+J_{\text{length}}$ (23)
@@
1:Initialize: Set $\mathbf{x}_{\text{ref}}(t)=\mathbf{x}_{0}(t)$, iteration counter $iter=0$, maximum trust region radius $\delta_{\text{max}}$
@@
$\min$ $J=J_{\text{smooth}}+J_{\text{length}}$ (25)
@@
9:     Compute consistency ratio $\rho$
$\rho=\frac{J(\mathbf{x}_{\text{ref}})-J(\mathbf{x})}{J_{\text{pred}}(\mathbf{x}_{\text{ref}})-J_{\text{pred}}(\mathbf{x})},$ (26)
10:     Increment iteration counter: $iter=iter+1$
11:until Convergence criterion is met
12:Output: Optimized trajectory

### v2

@@ 1 Introduction
In this study, we propose an enhanced Trust Region Sequential Convex Optimization algorithm tailored specifically for trajectory planning of multi-drone thermal screening systems in complex urban environments. Our contributions are threefold: First, we introduce an improved convexification method that effectively addresses the nonlinear and non-convex constraints characteristic of urban drone operations, resulting in enhanced convergence stability and computational efficiency. Second, the integration of adaptive trust-region filtering for constraint reduction and higher-order soft trust-region integration ensures robust trajectory refinement, significantly reducing computation time while maintaining path optimality and safety constraints. Finally, we validate the effectiveness and superiority of our algorithm through comprehensive simulations in realistic urban scenarios, demonstrating marked improvements in thermal screening coverage, trajectory smoothness, and overall system performance compared to conventional convex optimization methods. This research thus represents a significant advancement toward practical and scalable multi-drone thermal screening solutions in urban settings.
@@ 3.1 Introduction to the Trust Region Sequential Convex Optimization Algorithm
3.2 Enhancements for TR-SCO
3.2.1 Convexification
@@
To address the complex constraints and dynamic requirements of multi-drone thermal screening in urban environments, we introduce two key enhancements to the traditional TR-SCO framework, integrating insights from trust-region filtering and higher-order optimization strategies.
3.2.2 Adaptive Trust-Region Filtering for Constraint Reduction
Inspired by the trust-region filtered SCP (TRF-SCP) approach, we develop a dynamic constraint filtering mechanism to reduce computational complexity. Urban environments impose numerous collision-avoidance constraints (e.g., building obstacles and inter-drone spacing), but many constraints are inactive during specific flight segments. The filtering mechanism operates as follows: 1)Obstacle Avoidance Filter: For each drone’s reference trajectory $x_{k}^{r}(t)$, we define a trust region
$\|x_{k}(t)-x_{k}^{r}(t)\|_{2}\leq\delta$ (20)
and check if obstacles lie within this region. Obstacles outside the trust region are excluded from the current subproblem. 2)Inter-Drone Collision Filter: Similarly, we filter inter-drone collision constraints by checking if the distance between any two drones $\|x_{k}(t)-x_{m}(t)\|_{2}$ falls within the trust region. Only constraints that intersect the trust region are retained, significantly reducing the number of active constraints.
This filtering strategy reduces the convex subproblem size, enabling real-time optimization for large drone fleets. To balance exploration and convergence stability, we introduce an adaptive trust-region updating mechanism. The radius $\delta$ is adjusted iteratively based on constraint violation. If the constraint violation $\ell(X^{q})$ exceeds a threshold $\varepsilon_{\rho}$, the radius expands
$\delta^{q+1}=c_{2}\cdot\delta^{q},c_{2}>1$ (21)
to broaden the search space and resolve infeasibility. If $\ell(X^{q})\leq\varepsilon_{\rho}$, the radius contracts
$\delta^{q+1}=c_{1}\cdot\delta^{q},0<c_{1}<1$ (22)
to refine the solution and reduce linearization errors.
This mechanism dynamically eliminates inactive collision-avoidance constraints that lie outside the trust region, reducing the convex subproblem size. By focusing only on active constraints, the algorithm achieves faster computation compared to traditional SCP methods, making it feasible for real-time optimization in large drone fleets. This is critical for urban scenarios where dynamic obstacle avoidance and rapid trajectory updates are essential.
3.2.3 Higher-Order Soft Trust-Region Integration
To enhance solution optimality, we design a higher-order soft trust region, transforming the trust-region constraint from a first-order penalty to a higher-order term (e.g., $\delta^{p}$, $p>1$) in the objective function. This modification: 1)Guarantees KKT optimality: Higher-order terms (e.g., $p=2$) ensure that the converged solution satisfies the Karush–Kuhn–Tucker (KKT) conditions, whereas first-order terms may lead to suboptimal solutions. 2)Suppresses oscillation: The higher-order penalty penalizes large trust-region radii more strongly, stabilizing the iteration process and reducing fluctuations in trajectory smoothness.
The objective function is adjusted to:
$\min J=J_{\text{smooth}}+J_{\text{length}}+w_{\delta}\cdot\delta^{p}$ (23)
where $w_{\delta}$ is the penalty weight and $p\geq 2$ is the order. This ensures that thermal screening trajectories are both optimal and smooth, critical for maintaining infrared imaging accuracy.
Higher-order penalties enforce smaller trust-region radii near convergence, reducing jerk and improving thermal imaging stability. This is critical for maintaining infrared camera focus, as abrupt movements can introduce measurement errors.
@@ 3.3 Mathematical Formulation and Optimization Approach
$\min$ $J=J_{\text{smooth}}+J_{\text{length}}+w_{\delta}\cdot\delta^{p}$ (24)
@@
1:Initialize: Set $\mathbf{x}_{\text{ref}}(t)=\mathbf{x}_{0}(t)$, iteration counter $iter=0$
@@
$\min$ $J=J_{\text{smooth}}+J_{\text{length}}+w_{\delta}\cdot\delta^{p}$ (26)
@@
$\begin{cases}\delta^{q+1}=c_{2}\delta^{q},&c_{2}>1,\quad\text{if }\ell(X^{q})>\varepsilon_{\rho},\\ \delta^{q+1}=c_{1}\delta^{q},&0<c_{1}<1,\quad\text{if }\ell(X^{q})\leq\varepsilon_{\rho}\end{cases}$ (27)
9:     Increment iteration counter: $iter=iter+1$
10:until Convergence criterion is met
11:Output: Optimized trajectory